OpenStudy (anonymous):

3x - 3 = -18 x = -7 x = -5 x = -3 x = 5

OpenStudy (anonymous):

x=-5

OpenStudy (dmezzullo):

Show work plz

OpenStudy (anonymous):

3x-3=-18 +3 +3 ____________ 3x/3x=-15/3 x=-5

OpenStudy (nurali):

3x - 3 = -18 3x=-18+3 3x=-15 x=-15/3 x=-5

OpenStudy (luigi0210):

\[3x-3=-18\] Add 3\[3x=-15\] Divide\[x=-5\]

OpenStudy (nurali):

Anytime.